# Reverberation mapping of high-mass and high-redshift quasars using gravitational time delays

### v1

@@ 3.3 The lensed fraction
where is the lens density function and taken here to simply be the halo mass function (see eq. 3), is the redshift of the lens, is the angular diameter distance to it, and is the fraction of the critical area () in which sources would be multiply imaged (for example, for a SIS lens, ). The expression is the differential co-moving volume111Note that in some work the differential co-moving volume is defined such that it contains the factor seen in eq. 2., with where .
@@
where is the magnification distribution of lensed sources, is the survey flux limit, and is the cumulative luminosity function, i.e., the number of background sources brighter than .
@@
As is also indicated in Yue et al. (2022), out of all sources with apparent luminosity brighter than , the lensed fraction is given by (Mason et al., 2015a):
@@
After calculating the expected number of quasars to be detected as explained in the previous section, the number is multiplied by the lensed fraction for each redshift bin. Eventually, a total lensed fraction is calculated. We perform the calculation for three different regimes - quasars at all redshifts and luminosity ranges, high luminosity quasars (i.e. which corresponds to erg s-1) and high redshift quasars, namely . Last, out of the high redshift candidates we aim to obtain the number of quasars for which the notions we highlight here might be particularly useful, i.e, those with long lag times. We follow the same procedure, only this time for each redshift bin we find the lower limit for the luminosity such that the expected time lag of the C iv line is at least, e.g., a 100 days. As can be seen in Fig.1, the expected time lags for the emission line will be even longer.
@@ 4 Results
While it is encouraging to see that many AGN multiply imaged by galaxy clusters could be detected across the sky, it is also insightful to examine which fraction of this lensed population has a large expected time lag, such that the gain from lensing would be most significant. Columns 5 and 6 in Table 1 list the expected number of multiply imaged AGN that are expected to have lag times of, as an example, and days. For Euclid, we expect multiply imaged AGN with days observed time lags and with days; and with the Roman space telescope; and and with JWST, respectively, across the sky. This is about 1-2 orders of magnitude lower than the total number of high-redshift AGN that are multiply imaged by galaxy clusters, but shows that there exists a lensed population in the sky which can gain substantially from RM campaigns to probe the R-L relation at high-redshift, as we also discuss further below (section 5).
@@ 5 Discussion
To examine the feasibility of the idea proposed in this work, in the first part of this work we have shown the distribution of time delays by typical lensing clusters and compared it to the time lag distribution of quasars across their mass range and at various redshifts. After showing that indeed galaxy clusters supply useful time delays that can easily be as large as the the time lags expected for highest luminosity and highest redshift AGN, in the second part of this work we then estimated the number of AGN that are expected to be multiply imaged by galaxy clusters and observable with different observatories of interest.
@@
Before concluding the work and while perhaps quite intuitive, we now aim to provide a more visual demonstration of the idea, using an idealized case. In particular, we consider an illustrative scenario in which the expected time lag is equal to the time delay between two images of a quasar. We end by referring to a couple of recent lensed examples for which such RM observational campaigns may be feasible using existing facilities.
@@
To simulate a typical continuum light curve of a quasar, we use mock light curves of the quasar HE0435-1223 generated by H0LiCOW collaboration and available online (Wong et al., 2017; Chen et al., 2019) . We simulate the emission-line light curve by replicating the continuum light curve shifted by a certain time lag and add for illustration purposes, a Gaussian noise with a standard deviation in flux.
@@
In Fig. 3 we show the mock light curves for two lensed images A and B for a SL time delay of 1000 days, where it is in this case equal to the observed time lag. An observed BLR size of a 1000 days corresponds to a physical BLR size of 1000 light days. At a redshift of , for example, this corresponds to a SMBH luminosity of erg , according to the local scaling relations. For each image we show the same time-frame of 1200 days in the observer frame, while the duration was arbitrarily chosen for illustration purposes. For each image separately, similar to the case in which the quasar were unlensed, the pattern seen in the first 1000 days in continuum is seen repeating itself in emission lines a 1000 days later. In practice, to securely identify such an offset it would require at least several times this timeframe, hence order a decade or more of continuous monitoring. However, due to the TD being similar to the time lag, the same changes in the continuum seen in the trailing image are seen at the same time in emission lines in the leading image, essentially saving the years-long wait time to observe the lagging emission line changes. Suppose that in order to detect the year time lag a monitoring campaign spanning 15 years were needed for an unlensed source to accumulate enough data points and timespan (the true value depends on a variety of factors that we do not go into here). Investing 15 years of observing time would be looked upon as an uneasy and perhaps risky investment given that the output is uncertain. In contrast, imagine a lensed quasar such that the source is multiply imaged with at least three images, which - more often than not - would be the case for wide separation lensed quasars as we consider here (e.g., Sharon et al. 2017; Williams et al. 2021). In such a case the investment becomes much more reasonable: each observation triples the number of data-points compared to the unlensed case, and more importantly, assuming that then the corresponding emission line changes can be observed (even if not immediately identified) already in the first few batches of monitoring. Furthermore, each image would then be magnified, which would also ease the observational follow-up.
@@
As another example, in the case of a smaller BLR size, causing a shorter time lag than the SL time delay, the pattern seen in the first days of observations in continuum would already be seen repeated in the emission lines after a short time lag. This is similar to the way it would appear for an unlensed quasar, where as for the lensed case we mainly gain magnification and additional sampling from the different multiple images, making the reverberation mapping more efficient overall. Also in the opposite scenario where the SL time delay is much shorter than the expected time lag, while we do not gain much in terms of overall wait time, we gain in several other aspects as above. First, we obtain additional points in the light curve with the same observation, compared to the unlensed case (an advantage that increases even further when more than two quasars images are seen, as is often the case in cluster lensing). Second, the images are magnified, which facilitates their follow-up. Third, while of singly imaged (or unlensed) objects can only be observed for several months a year, when they are away from the Sun’s direction, a multiply-imaged quasar may allow us to fill gaps in the light curves also in those otherwise down times.
It should however be noted that even if TDs are used to facilitate RM, the monitoring time should still be sufficiently significant so as not to bias the measurements towards the available TDs.
Recently, a successful reverberation mapping was performed on the multiply imaged quasar, SDSS J2222+2745, at (Williams et al., 2021), demonstrating the usefulness of multiply imaged quasars in that respect. A time lag of restframe days was significantly detected for the C iv emission line, whereas the time delays to the two other images of the quasar are about and days (see also Sharon et al. 2017). Hence, one time delay is of the order of the time lag, similar to first case discussed above (but with a shorter time lag than what we consider here, which is aimed for the high-redshift and high luminosity regimes), and the second time delay is much longer than the time lag, similar to second case discussed above.
@@
Note that throughout we assume that any object detected by a certain observatory could also be followed up spectroscopically (by the same or other observatories). In addition, given that we only wish to demonstrate the basic idea, we only show an idealized case. We also neglect here other sources of noise such as local microlensing for some of the images, which is known to often be present and cause its own variations to the light curve for some of the components (e.g., Eigenbrod et al., 2008; Mediavilla et al., 2015; Vernardos et al., 2024) – depending on the mass of the point mass and size of the emitting source – although even in such cases the effects can be remedied (see e.g. Williams et al. (2021)). Similarly, we do not take into account so-called “changing-look” AGN (e.g. Ricci & Trakhtenbrot, 2023), for which some of the images may not show, for example, broad-line emission, while the other images might (in fact, lensing TDs may be a route for detecting lensed AGN of this type).
@@
One should also bear in mind that in reality the TD is not necessarily known from the lens model to the required accuracy to remove it from the fit altogether. However, it might be first determined from the temporal shift between the continua of the two images before determining the time lag. This may improve the error on the TDs compared to the errors from the lens model’s TD, which would otherwise propagate into the R-L relation as well.
@@
It is also worth highlighting a couple of interesting recent objects which may become useful targets for RM of lensed AGN (other examples naturally exist, too). First we note the triply-imaged Little Red Dot AGN reported by Furtak et al. (2023b) at . This object is currently the highest-redshift multiply imaged AGN. It has according to their lens model a TD of 3 years, roughly, between the two closest images, and a TD of about 19 years between the first of these two and the third, more distant image, which is first to arrive. The NIRSpec spectrum of the source was published in Furtak et al. (2024). Based on its luminosity, with a few hours per visit we could detect both the continuum and the H line (no C iv lines were detected in the object, despite the very deep integration time). One can thus envision that a NIRSpec/JWST program, for example, targeting the three images simultaneously, could be used to measure both the continuum and emission lines in each epoch. From the width of H, Furtak et al. (2024) derived a BH mass of 4 M⊙, which, according to the scaling relations by Kaspi et al. (2021), corresponds to an expected or observed time lag of a couple dozen days. This time lag is relatively short (corresponding to the second case discussed above), and while its measurement could in principle also be done were it not lensed, its measurement would be facilitated by measuring the three images simultaneously. While the expected time lag for this object does not appear to be very long (so that the gain in wait time is not maximized), it could supply a unique example for probing the RM relation at high-redshift.
@@
A second, recent example is the large separation quasar COOL J0335-1927 (Napier et al., 2023a), for which time delays of days and days between two pairs of its images were predicted from the lens model, although with quite substantial uncertainty. To estimate the expected time lag, we adopt the typical -band luminosity seen in Fig. 4 in their work, and translate it to an absolute UV magnitude of AB, which, based on existing scaling relations (Kaspi et al., 2021), corresponds to a an observed time lag of days. The object therefore might benefit significantly from such a campaign. For more discussion on the latter object, see also Napier et al. (2023a).
@@
We thus conclude that designated observations of multiply imaged quasars can open the door for RM studies in a much more efficient manner, as was already demonstrated at lower-redshifts Williams et al. (2021). For the most massive and/or high-redshift quasars for which the time lags may be otherwise too long, this might possibly be the only, feasible way.
@@ 6 Summary
In this work we discuss the possibility of using lensing time delays to facilitate the RM of high-luminosity and high-redshift quasars, for which the radius-luminosity relation is currently, poorly constrained. The expected time lags for very-massive or high-redshift quasars can reach several years and would likely require much longer monitoring times than can be comfortably accommodated by typical observing campaigns. Multiply imaged AGN may, however, offer a unique opportunity to explore the radius-luminosity relation at these ends. Not only do strongly lensed quasars comprise several magnified images which boost the efficiency of the light curve sampling, but also, the time delay between the quasar’s multiple images can save the long wait times that are typically required. For example, if the time delay is roughly of the order of the expected time lag, changes in the emission lines in a leading image can be observed around the same time as the changes in the continuum in a trailing image, thus allowing a much reduced wait times and overall monitoring times-spans.
We start by showing the typical distribution of BH time lags compared to the lensing TD distribution from cluster lenses and show that a strong overlap exists. We then estimate the number of multiply imaged quasars that may be detectable in the near future with Euclid, Roman, or JWST. We find that indeed, a fair amount of multiply imaged high-redshift quasars, out of which some would experience a significantly long observed time lag, should be detectable. Using the time delays between the multiple images, RM of these possible future candidates could be obtained in a reasonable few-year monitoring timescale, probing the poorly known R-L relation in the early Universe.
@@ Acknowledgements
M.G would like to thank Yuichi Harikane for a useful discussion. M.G is thankful for the hospitality of the SURF program and cosmic DAWN center where much of this work has been carried out. The BGU group acknowledges support by grant No. 2020750 from the United States-Israel Binational Science Foundation (BSF) and grant No. 2109066 from the United States National Science Foundation (NSF); by the Israel Science Foundation Grant No. 864/23; and by the Ministry of Science & Technology, Israel. D.C. acknowledges support by grants from the German Research Foundation (HA3555-14/1, CH71-34-3) and the Israeli Science Foundation (2398/19, 1650/23). This research made use of Astropy,222http://www.astropy.org a community-developed core Python package for Astronomy (Astropy Collaboration et al., 2013; Price-Whelan et al., 2018) as well as the packages cluster_toolkit, CLASS (Blas et al., 2011), NumPy (van der Walt et al., 2011), SciPy (Virtanen et al., 2020) and the astronomy library for MATLAB (Ofek, 2014). The Matplotlib package (Hunter, 2007) was used to create some of the figures in this work.

### v2

@@ 3.3 The lensed fraction
where is the lens density function and taken here to simply be the halo mass function (see eq. 3), is the redshift of the lens, is the angular diameter distance to it, and is the fraction of the critical area () in which sources would be multiply imaged (for example, for a SIS lens, ). The expression is the differential co-moving volume222Note that in some work the differential co-moving volume is defined such that it contains the factor seen in eq. 2., with where .
@@
where is the magnification distribution of lensed sources, is the survey flux limit, and is the cumulative luminosity function, i.e., the number of background sources more luminous than .
@@
As is also indicated in Yue et al. (2022), out of all sources with apparent luminosity greater than , the lensed fraction is given by (Mason et al., 2015a):
@@
Note. — In every column, we also specify within the brackets the total number of quasars expected to be detected; Column 1: The assumed QLF. The detection limit is calculated assuming a blue selection unless otherwise specified; Column 2: The estimated number of lensed quasars at all redshifts and all luminosity range; Column 3: The estimated number of lensed quasars at all redshifts and high luminosity range; Column 4: The estimated number of lensed quasars at high redshifts; Columns 5-6: The estimated number of lensed quasars with luminosities and redshift that would result in a C iv observed RM time lag of above 50 and 100 days;
After calculating the expected number of quasars to be detected as explained in the previous section, the number is multiplied by the lensed fraction for each redshift bin. Eventually, a total lensed fraction is calculated. We perform the calculation for three different regimes - quasars at all redshifts and luminosity ranges, high luminosity quasars (i.e. which corresponds to erg s-1) and high redshift quasars, namely . Last, out of the high redshift candidates we aim to obtain the number of quasars for which the notions we highlight here might be particularly useful, i.e, those with long lag times. We follow the same procedure, only this time for each redshift bin we find the lower limit for the luminosity such that the expected RM time lag of the C iv line is at least, e.g., a 100 days. As can be seen in Fig.1, the expected RM time lags for the emission line will be even longer.
@@ 4 Results
While it is encouraging to see that many AGN multiply imaged by galaxy clusters could be detected across the sky, it is also insightful to examine which fraction of this lensed population has a large expected RM time lag, such that the gain from lensing would be most significant. Columns 5 and 6 in Table 1 list the expected number of multiply imaged AGN that are expected to have lag times of, as an example, and days. For Euclid, we expect multiply imaged AGN with days observed RM time lags and with days; and with the Roman space telescope; and and with JWST, respectively, across the sky. This is about 1-2 orders of magnitude lower than the total number of high-redshift AGN that are multiply imaged by galaxy clusters. All in all, our calculations show that there exists a large lensed population in the sky which can gain substantially from RM campaigns to probe the R-L relation at high-redshift. This may be particularly beneficial for those with longer RM lag times, as we also discuss further below (section 5).
@@ 5 Discussion
To examine the feasibility of the idea proposed in this work, in the first part of this work we have shown the distribution of TDs by typical lensing clusters and compared it to the RM time lag distribution of quasars across their mass range and at various redshifts. After showing that indeed galaxy clusters supply useful TDs that can easily be as large as the RM time lags expected for highest luminosity and highest redshift AGN, in the second part of this work we then estimated the number of AGN that are expected to be multiply imaged by galaxy clusters and observable with different observatories of interest. We find that for all three observatories we consider here, namely JWST, Euclid, and Roman Space Telescope, several hundreds to several tens of thousands of lensed AGN should be discovered by these observatories to typical depths (see Table 1).
@@
Before concluding the work and while perhaps quite intuitive, we now aim to provide a more visual demonstration of the idea, using an idealized case. In particular, we consider an illustrative scenario in which the expected RM time lag is equal to the TD between two images of a quasar. We then discuss some of the observational challenges, that are not taken into account in the idealized example. We end by referring to a couple of recent lensed examples for which such RM observational campaigns may be feasible using existing facilities.
5.1 Idealized light curve example
@@
To simulate a typical continuum light curve of a quasar, we use mock light curves of the quasar HE0435-1223 generated by H0LiCOW collaboration and available online (Wong et al., 2017; Chen et al., 2019) . We simulate the emission-line light curve by replicating the continuum light curve shifted by a certain RM time lag and add for illustration purposes, a Gaussian noise with a standard deviation in flux.
@@
In Fig. 4 we show the mock light curves for two lensed images A and B for a SL TD of 1000 days, where it is in this case equal to the observed RM time lag. An observed BLR size of a 1000 days corresponds to a physical BLR size of 1000 light days. At a redshift of , for example, this corresponds to a SMBH luminosity of erg , according to the local scaling relations. For each image we show the same time-frame of 1200 days in the observer frame, while the duration was arbitrarily chosen for illustration purposes. For each image separately, similar to the case in which the quasar were unlensed, the pattern seen in the first 1000 days in continuum is seen repeating itself in emission lines a 1000 days later. In practice, to securely identify such an offset it would require at least several times this timeframe, hence order a decade or more of continuous monitoring. However, due to the TD being similar to the RM time lag, the same changes in the continuum seen in the trailing image are seen at the same time in emission lines in the leading image, essentially saving the years-long wait time to observe the lagging emission line changes. Suppose that in order to detect the year RM time lag a monitoring campaign spanning 15 years were needed for an unlensed source to accumulate enough data points and timespan (the true value depends on a variety of factors that we do not go into here). Investing 15 years of observing time would be looked upon as an uneasy and perhaps risky investment given that the output is uncertain. In contrast, imagine a lensed quasar such that the source is multiply imaged with at least three images, which - more often than not - would be the case for wide separation lensed quasars as we consider here (e.g.,Sharon et al. 2017; Williams et al. 2021). In such a case the investment becomes much more reasonable: each observation triples the number of data-points compared to the unlensed case, and more importantly, assuming that then the corresponding emission line changes can be observed (even if not immediately identified) already in the first few batches of monitoring. Furthermore, each image would then be magnified, which would also ease the observational follow-up.
@@
As another example, in the case of a smaller BLR size, causing a shorter RM time lag than the SL TD, the pattern seen in the first days of observations in continuum would already be seen repeated in the emission lines after a short RM time lag. This is similar to the way it would appear for an unlensed quasar, where as for the lensed case we mainly gain magnification and additional sampling from the different multiple images, making the reverberation mapping more efficient overall. Also in the opposite scenario where the SL TD is much shorter than the expected RM time lag, while we do not gain much in terms of overall wait time, we gain in several other aspects as above. First, we obtain additional points in the light curve with the same observation, compared to the unlensed case (an advantage that increases even further when more than two quasars images are seen, as is often the case in cluster lensing). Second, the images are magnified, which facilitates their follow-up. Third, while singly imaged (or unlensed) objects can only be observed for several months a year, when they are away from the Sun’s direction, a multiply-imaged quasar may allow us to fill gaps in the light curves also in those otherwise down times.
5.2 Challenges and caveats
After demonstrating how TDs can be exploited for RM in an idealized case, we now refer to some challenges that might arise in actual RM campaigns. First, it should be noted that throughout we assume that any object detected by a certain observatory could also be followed up spectroscopically (by the same or other observatories). In practice, follow-up may be of course more challenging, especially for those objects with larger TDs or those only reachable from space, yet in any case for all of these objects there is a gain from lensing compared to the unlensed case.
In addition, we note that even if TDs are used to facilitate RM, the monitoring time should still be sufficiently significant so as not to bias the measurements towards the available TDs. A substantial monitoring time would be needed to also account for the fact that in practice, emission-line light curves tend to be smoother compared to the continuum light curves, adding to challenges involved in assigning specific changes in the emission lines to the correct continuum changes.
@@ 5.2 Challenges and caveats
We also neglect here other sources of noise such as local microlensing for some of the images, which is known to often be present and cause its own variations to the light curve for some of the components (e.g., Eigenbrod et al., 2008; Mediavilla et al., 2015; Vernardos et al., 2024) – depending on the mass of the point mass and size of the emitting source – although even in such cases the effects can be remedied (see e.g. Williams et al. (2021)). Similarly, we do not take into account so-called “changing-look” AGN (e.g. Ricci & Trakhtenbrot, 2023), for which some of the images may not show, for example, broad-line emission, while the other images might. This population of AGN seems to be extremely rare, with only about 150 discovered to date (Wang et al., 2023) and would most likely not effect our calculations. On the other hand, lensing TDs may be a route for detecting lensed AGN of this type.
@@
One should also bear in mind that in reality the TD is not necessarily known from the lens model to the required accuracy to remove it from the fit altogether. However, it might be first determined from the temporal shift between the continua of the two images before determining the RM time lag. This may improve the error on the TDs compared to the errors from the lens model’s TD, which would otherwise propagate into the R-L relation as well.
5.3 Future prospects
Recently, a successful reverberation mapping was performed on the multiply imaged quasar, SDSS J2222+2745, at (Williams et al., 2021), demonstrating the usefulness of multiply imaged quasars in that respect. A RM time lag of restframe days was significantly detected for the C iv emission line, whereas the TDs to the two other images of the quasar are about and days (see also Sharon et al. 2017). Hence, one TD is of the order of the RM time lag, similar to first case discussed in 5.1 (but with a shorter RM time lag than what we consider here, which is aimed for the high-redshift and high luminosity regimes), and the second TD is much longer than the RM time lag, similar to second case discussed in 5.1.
@@ 5.3 Future prospects
It is now worth highlighting also a couple of other, interesting recent objects which may become useful targets for RM of lensed AGN (other examples naturally exist, too). First we note the triply-imaged Little Red Dot AGN reported by Furtak et al. (2023b) at . This object is currently the highest-redshift multiply imaged AGN. It has according to their lens model a TD of 3 years, roughly, between the two closest images, and a TD of about 19 years between the first of these two and the third, more distant image, which is first to arrive. The NIRSpec spectrum of the source was published in Furtak et al. (2024). Based on its luminosity, with a few hours per visit we could detect both the continuum and the H line (no C iv lines were detected in the object, despite the very deep integration time). One can thus envision that a NIRSpec/JWST program, for example, targeting the three images simultaneously, could be used to measure both the continuum and emission lines in each epoch. From the width of H, Furtak et al. (2024) derived a BH mass of 4 M⊙, which, according to the scaling relations by Kaspi et al. (2021), corresponds to an expected or observed RM time lag of a couple dozen days. This RM time lag is relatively short (corresponding to the second case discussed in 5.1), and while its measurement could in principle also be done were it not lensed, its measurement would be facilitated by measuring the three images simultaneously. While the expected RM time lag for this object does not appear to be very long (so that the gain in wait time is not maximized), it could supply a unique example for probing the RM relation at high-redshift.
@@
A second, recent example is the large separation quasar COOL J0335-1927 (Napier et al., 2023a), for which TDs of days and days between two pairs of its images were predicted from the lens model, although with quite substantial uncertainty. To estimate the expected RM time lag, we adopt the typical -band luminosity seen in Fig. 4 in their work, and translate it to an absolute UV magnitude of AB, which, based on existing scaling relations (Kaspi et al., 2021), corresponds to a an observed RM time lag of days. The object therefore might benefit significantly from such a campaign. For more discussion on the latter object, see also Napier et al. (2023a).
@@
We thus conclude that designated observations of multiply imaged quasars can open the door for RM studies in a much more efficient manner, as was already demonstrated at lower-redshifts (Williams et al., 2021). For the most massive and/or high-redshift quasars for which the RM time lags may be otherwise too long, this might possibly be the only, feasible way.
@@ 6 Summary
In this work we discuss the possibility of using lensing TDs to facilitate the RM of high-luminosity and high-redshift quasars, for which the radius-luminosity relation is currently, poorly constrained. The expected RM time lags for very-massive or high-redshift quasars can reach several years and would likely require much longer monitoring times than can be comfortably accommodated by typical observing campaigns. Multiply imaged AGN may, however, offer a unique opportunity to explore the radius-luminosity relation at these ends. Not only do strongly lensed quasars comprise several magnified images which boost the efficiency of the light curve sampling, but also, the TD between the quasar’s multiple images can save the long wait times that are typically required. For example, if the TD is roughly of the order of the expected RM time lag, changes in the emission lines in a leading image can be observed around the same time as the changes in the continuum in a trailing image, thus allowing a much reduced wait times and overall monitoring times-spans.
We start by showing the typical distribution of BH RM time lags compared to the lensing TD distribution from cluster lenses and show that a strong overlap exists. We then estimate the number of multiply imaged quasars that may be detectable in the near future with Euclid, Roman, or JWST. We find that indeed, a fair amount of multiply imaged high-redshift quasars, out of which some would experience a significantly long observed RM time lag, should be detectable. Using the TDs between the multiple images, RM of these possible future candidates could be obtained in a reasonable few-year monitoring timescale, probing the poorly known R-L relation in the early Universe.
@@ Acknowledgements
M.G would like to thank Yuichi Harikane for a useful discussion. M.G is thankful for the hospitality of the SURF program and cosmic DAWN center where much of this work has been carried out. The BGU group acknowledges support by grant No. 2020750 from the United States-Israel Binational Science Foundation (BSF) and grant No. 2109066 from the United States National Science Foundation (NSF); by the Israel Science Foundation Grant No. 864/23; and by the Ministry of Science & Technology, Israel. D.C. acknowledges support by grants from the German Research Foundation (HA3555-14/1, CH71-34-3) and the Israeli Science Foundation (2398/19, 1650/23). This research made use of Astropy,333http://www.astropy.org a community-developed core Python package for Astronomy (Astropy Collaboration et al., 2013; Price-Whelan et al., 2018) as well as the packages cluster_toolkit, CLASS (Blas et al., 2011), NumPy (van der Walt et al., 2011), SciPy (Virtanen et al., 2020) and the astronomy library for MATLAB (Ofek, 2014). The Matplotlib package (Hunter, 2007) was used to create some of the figures in this work.